### v3

@@ Abstract
While computer vision has achieved tremendous success with multimodal encoding and direct textual interaction with images via chat-based large language models, similar advancements in medical imaging AI—particularly in 3D imaging—have been limited due to the scarcity of comprehensive datasets. To address this critical gap, we introduce CT-RATE, the first dataset that pairs 3D medical images with corresponding textual reports. CT-RATE comprises 25,692 non-contrast 3D chest CT scans from 21,304 unique patients. Through various reconstructions, these scans are expanded to 50,188 volumes, totaling over 14.3 million 2D slices. Each scan is accompanied by its corresponding radiology report. Leveraging CT-RATE, we develop CT-CLIP, a CT-focused contrastive language-image pretraining framework designed for broad applications without the need for task-specific training. We demonstrate how CT-CLIP can be used in two tasks: multi-abnormality detection and case retrieval. Remarkably, in multi-abnormality detection, CT-CLIP outperforms state-of-the-art fully supervised models across all key metrics, effectively eliminating the need for manual annotation. In case retrieval, it efficiently retrieves relevant cases using either image or textual queries, thereby enhancing knowledge dissemination. By combining CT-CLIP’s vision encoder with a pretrained large language model, we create CT-CHAT, a vision-language foundational chat model for 3D chest CT volumes. Finetuned on over 2.7 million question-answer pairs derived from the CT-RATE dataset, CT-CHAT surpasses other multimodal AI assistants, underscoring the necessity for specialized methods in 3D medical imaging. Collectively, the open-source release of CT-RATE, CT-CLIP, and CT-CHAT not only addresses critical challenges in 3D medical imaging, but also lays the groundwork for future innovations in medical AI and improved patient care.
Introduction
@@
Three-dimensional (3D) medical imaging, such as computed tomography (CT) and magnetic resonance imaging (MRI), provides a more detailed and comprehensive view of a patient’s condition compared to two-dimensional (2D) imaging [1]. Among these, CT stands out due to its speed and widespread availability, making it the preferred modality for diagnosing various abnormalities [2]. As global demand for CT scans continues to grow, the need for rapid and accurate diagnostic results becomes more pressing. However, the manual generation of reports from these complex images is time-consuming and prone to errors, underscoring the urgent need for automation [3]. Artificial intelligence (AI) has the potential to address this challenge by efficiently identifying pathologies, reducing the workload on radiologists, and potentially saving lives in situations where immediate radiological expertise is not available.
@@
AI has already begun to revolutionize the interpretation of medical images [4, 5, 6, 7, 8]. Recent advances in automated diagnosis, classification, and segmentation have been driven by numerous publicly available datasets [9, 10, 11, 12, 13]. Beyond supervised methods, general AI systems have been developed using open-source image-text pairs in self-supervised learning, particularly in 2D imaging [14]. These versatile AI foundation models integrate comprehensive natural language annotations into the learning process without requiring task-specific training, thereby eliminating the need for explicit manual image annotations. Unconstrained by predefined classes or tasks during training, they enhance adaptability and have been successfully applied to tasks such as disease classification [15], case retrieval [16], and interactive diagnostics via chat interfaces [14]. Despite these significant advancements in 2D imaging, the application of advanced AI techniques to 3D imaging remains underexplored. A major challenge is the lack of open-source 3D medical imaging datasets paired with textual reports, which require extensive resources, specialized knowledge, and technical expertise to create [17, 18, 19]. Currently, the largest 3D medical imaging datasets available contain only a few thousand cases and do not include paired textual data [20, 21, 22], limiting the development of advanced AI models. Moreover, 3D models face higher computational demands and have fewer available pretrained open-source models compared to their 2D counterparts [23, 24]. Previous studies have only begun to address these challenges; for example, one involving chest CT is limited by its supervised approach, restricting detection to predefined abnormalities and offering only a small portion of its dataset publicly without paired textual reports [20].
@@
To address these challenges, we present three related contributions: First, we introduce CT-RATE, a dataset that pairs 50,188 reconstructed 3D chest CT volumes (over 14.3 million 2D slices) from 25,692 scans of 21,304 patients with corresponding radiology reports. Each volume is annotated with 18 abnormalities, supporting extensive validation and fine-tuning. CT-RATE is fully open-source, making it the first public 3D medical imaging dataset with paired textual reports and the largest CT dataset with multi-abnormality labels. Second, leveraging CT-RATE, we develop CT-CLIP, a novel chest CT-focused contrastive language-image pretraining framework. CT-CLIP integrates extensive natural language annotations into the learning process, enabling the model to comprehend image-based semantic knowledge and perform various tasks without specific training. Our evaluation shows CT-CLIP’s ability to detect multiple abnormalities in a zero-shot setting [25], outperforming state-of-the-art fully supervised methods on three test sets from two different countries. These findings indicate that explicit labels are not essential for achieving high performance in 3D medical imaging interpretation, given the availability of corresponding radiology reports. Third, to extend CT-CLIP’s capabilities, we combine its vision encoder with a pretrained large language model to create CT-CHAT, a vision-language chat model for 3D chest CT volumes. CT-CHAT is fine-tuned with over 2.7 million question-answer pairs derived from the CT-RATE dataset. CT-CHAT significantly outperforms open-source conversational vision-language AI assistants, such as LLaVA 1.6 [26], LLaVA-Med [27], and CXR-LLaVA [28], underscoring the need for specialized tools in 3D medical imaging. We also benchmark CT-CHAT against state-of-the-art 3D medical report generation methods, including RadFM [29] and CT2Rep [30], using clinically meaningful accuracy metrics. On two independent test sets, CT-CHAT outperforms these baselines by a significant margin, demonstrating both its clinical relevance and technical efficacy in report generation.
In this article, we introduce pivotal contributions to medical AI. We address the scarcity of open-source 3D imaging datasets paired with textual reports by introducing CT-RATE. Then, we create CT-CLIP, a versatile foundation model that effectively integrates linguistic and visual data from CT-RATE via contrastive learning [31]. We demonstrate CT-CLIP’s remarkable adaptability and performance across a variety of tasks without task-specific training. Moreover, we develop a vision-language AI assistant, CT-CHAT, by finetuning a pretrained large language model through visual instruction using CT encodings created by CT-CLIP’s vision encoder. The open-source release of CT-RATE, CT-CLIP, and CT-CHAT addresses critical challenges in 3D medical imaging, such as dataset scarcity, model versatility, and learning paradigms, leading to groundbreaking research and substantial advancements in the field.
@@ Results
We curate and open-source CT-RATE, a dataset comprising 25,692 non-contrast 3D chest CT volumes and corresponding radiology text reports from 21,304 unique patients (Figure 1a). These 3D volumes, reconstructed using various methods to suit different window settings [32], total over 14.3 million 2D slices. Each CT exam is accompanied by a report that includes an impression and a findings section, authored by a radiologist, along with additional sections detailing patient information and scan technique. CT-RATE is divided into training and internal validation sets for experiments (see creating the CT-RATE dataset section of Methods). While CT-RATE is a single-center dataset, it captures substantial heterogeneity in CT scanners, acquisition protocols, and reconstruction techniques. To evaluate the generalization of our models (developed on CT-RATE) under distribution shifts, we incorporate two external validation sets from independent studies conducted in different countries [20, 33] (see adding external validation datasets to CT-RATE section of Methods). First, we utilize the public RAD-ChestCT dataset, which includes multi-abnormality labels, to assess model generalizability. Second, we perform a real-world validation by inviting an independent clinical team—uninvolved in model development—to download our open-source codebase, models, and user interface from the public repository and apply them to their own local, private dataset without any adaptation. The performance scores they obtain are presented in the following sections. Notably, none of the external validation data are used during development; all model development is conducted exclusively using CT-RATE.
@@
Figure 2 summarizes the CT-RATE dataset alongside our open-source external validation set, RAD-ChestCT. Our dataset spans a wide age range (18-102 years) (Figure 2a) and includes diverse pathologies primarily related to both heart and lung conditions (Figure 2b). The reports vary in length and are authored by numerous radiologists, reflecting a variety of writing styles (Figure 2c). The dataset also varies in CT slice numbers (Figure 2d), slice thickness, axial pixel (XY) spacing (Figure 2e), and includes data from different CT manufacturers and a balanced sex distribution (Figure 2f). This heterogeneity is crucial for developing robust models capable of handling out-of-distribution cases. Supplementary Table 1 provides detailed statistics on abnormalities, manufacturers, and sex distributions in the training and internal validation sets. In conclusion, CT-RATE does not represent a curated study dataset, but rather reflects real-world scenarios characterized by heterogeneity. This diversity ensures that the models trained on CT-RATE are well-suited for deployment in clinical settings, offering reliable performance across a broad range of conditions and imaging scenarios.
@@
Unlike other methods utilizing supervised machine learning for multi-abnormality classification of CT volumes, our study employs radiology reports, which provide a rich source of semantic knowledge. This significantly enhances the understanding of CT volumes and supports a wide range of potential zero-shot applications. Accordingly, we develop CT-CLIP, the first foundational model for chest CT-focused contrastive language-image pretraining, adapting a technique from a previous study [37]. We utilize both the impression and findings sections of radiology reports combined for CT-CLIP due to their superior performance compared to utilizing either one alone, as demonstrated by the abnormality classification scores in an ablation study (Supplementary Figure 1a). While the best performance is achieved training on both impression and findings together, using only the findings section performs better than using only the impression section, indicating that the findings section contains more clinically informative content.
@@
We compare our zero-shot approach to this supervised baseline and observe significantly higher performance metrics for our approach. On the internal validation set from CT-RATE, the zero-shot CT-CLIP model achieves a mean AUROC that is 0.102 higher and a mean F1 score that is 0.050 higher than the supervised baseline. On the first external validation set, RAD-ChestCT, CT-CLIP outperforms the baseline with a mean AUROC gain of 0.085 and an F1 score improvement of 0.073. Similarly, on the second external validation set from UPMC, CT-CLIP maintains strong performance with a 0.087 increase in mean AUROC and a 0.037 boost in F1 score.Results on the external validation sets demonstrate that the zero-shot CT-CLIP model performs well not only on similar data distributions but also under distribution shifts, indicating its strong generalization to external datasets. A two-sided unpaired permutation test, utilizing 10,000 permutations, highlights a statistically significant improvement $(p<0.05)$ in AUROC as a zero-shot classifier over the supervised (baseline) model, detailed in Supplementary Table 5.
Extended Data Figures 3, 4, 5 provide scores for all abnormalities on the CT-RATE internal, RAD-ChestCT external and UPMC external validation sets, respectively. Supplementary Figures 3, 7 show confusion matrices on the CT-RATE internal and RAD-ChestCT external validation sets, respectively. Notably, on the internal validation set, all abnormalities (18 out of 18) exhibit higher scores in our zero-shot approach compared to the supervised baseline in mean AUROC. In conclusion, CT-CLIP excels by just leveraging existing radiology text reports in electronic health records, thus avoiding the need for expensive manual annotations. Moreover, CT-CLIP’s zero-shot capability offers flexible predictions beyond predefined classes and adapts efficiently to changing needs post-training.
Since the availability of sufficiently large datasets is crucial for machine learning, especially for transformer models, which require more data than convolutional neural networks [40, 41], we explore the impact of dataset size on classification performance. We train CT-CLIP on different portions of the CT-RATE dataset, specifically 9.8% (equivalent to the open-source portion of the RAD-ChestCT dataset [20], currently the largest publicly available chest CT dataset), 20%, 40%, 60%, 80%, and 100%. Figure 4c reports the classification performance across different dataset sizes. Notably, a consistent trend emerges: larger dataset sizes correlate with improved scores. This finding highlights the role of dataset size in enhancing CT-CLIP’s zero-shot classification performance and underscores the value of our extensive CT-RATE dataset for computational research. It also offers a clear avenue for future improvements, suggesting that further expanding the dataset or incorporating more diverse data could lead to even better performance.
@@
To enhance CT-CLIP’s multi-abnormality detection performance, we explore a task-specific finetuning method, ClassFine, inspired by conventional linear probing, by adding a new classification layer to the pretrained vision transformer of CT-CLIP (Figure 3a). Supplementary Figure 1b indicates that training only the final layer may be sufficient for finetuning (see finetuning CT-CLIP for multi-abnormality detection section of Methods). While the ClassFine approach improves detection performance (Figure 4a), it restricts the model to predefined labels as any supervised classification method. Extended Data Figures 3, 4, 5 present abnormality scores, with confusion matrices in Supplementary Figures 5, 9 on the internal and external validation sets, respectively. To address the limitations of ClassFine, we develop the VocabFine approach, a novel open-vocabulary finetuning method inspired by WISE-FT [42] (see finetuning CT-CLIP for multi-abnormality detection section of Methods). VocabFine, like CT-CLIP’s zero-shot inference, is finetuned to improve label prediction accuracy (Figure 3c). Specifically, we use positive and negative prompts for abnormalities, calculate the logits for these volumes with pretrained CT-CLIP, and align these logits with the ground truth labels. While the finetuned model with VocabFine retains the open-vocabulary inference capability of CT-CLIP (Figure 3d), it also consistently outperforms ClassFine (Figure 4a), further emphasizing the value of our open vocabulary finetuning. Abnormality-specific improvements are detailed in Extended Data Figures 3, 4, 5 while Supplementary Figure 1b underscores the importance of training all weights in finetuning. Supplementary Figures 4, 8 provide confusion matrices on the internal and external validation sets, respectively. Notably, both ClassFine and VocabFine are trained solely on CT-RATE; thus, their strong generalization to external sets (RAD-ChestCT and UPMC) demonstrates the robustness of CT-CLIP’s pretrained representations under distribution shifts. This performance advantage across multiple test sets supports the effectiveness of our finetuning strategies. However, despite these promising results, the abnormality detection performance still falls short of clinical deployment thresholds.
@@
Beyond its superior performance, VocabFine retains the open-vocabulary capabilities of the original CT-CLIP model, enabling inference on abnormalities that were not explicitly included during fine-tuning. This makes the model inherently more flexible and clinically useful in real-world scenarios where emerging or rare findings may need to be recognized without retraining. Additionally, unlike ClassFine—which only adds a classification head on top of the vision encoder and is therefore restricted to vision-only downstream tasks—VocabFine preserves and fine-tunes both the vision and text encoders. This dual-modality design allows it to generalize to a wider range of tasks, such as text-to-image retrieval and multimodal question answering, which rely on aligned visual and linguistic representations. In summary, our VocabFine approach not only improves multi-abnormality detection accuracy but also maintains the foundational model’s versatility and adaptability across diverse downstream applications.
@@
Beyond zero-shot detection, CT-CLIP serves as a powerful tool for image-to-image retrieval [43], identifying the relevant volumes from a query volume by comparing cosine similarities between the embeddings (Figure 5a). We test on both CT-RATE internal and RAD-ChestCT external validation sets using the vision transformer of CT-CLIP, finetuned models, and the supervised baseline. We utilize the Mean Average Precision at K (MAP@K) metric, where K represents the number of returned images (see benchmarking the CT-CLIP models for CT volume retrieval section of Methods). As K increases, MAP@K scores typically decrease due to less accurate matches. Results, as in Figure 5c, show comparable performance among methods, with all significantly outperforming random retrieval. Although the external validation set from RAD-ChestCT has higher MAP@K scores, the fold increase over random retrieval is lower (three times on the external vs. six times on the internal set), due to differing abnormality distributions. This suggests that examining fold change offers a more accurate comparison. Finetuning does not significantly alter the vision transformer’s weights, as CT-CLIP and the finetuned models perform similarly across MAP@K metrics.
@@
CT-CHAT’s architecture leverages the strengths of CT-CLIP by using its pretrained vision transformer to encode 3D chest CT volumes. These encodings are then passed to a large language model (LLM) through a multimodal projector, as shown in Figure 6a. We perform an ablation study evaluating CT-CHAT’s performance with 4 different state-of-the-art LLMs, as detailed in Figure 6d. We select 70B Meta Llama 3.1 [45], outperforming others in all key metrics, for the optimal configuration of CT-CHAT.
@@
Given the absence of generalist models for 3D imaging, we evaluate CT-CHAT against several state-of-the-art open-source vision-language AI models, including LLaVA 1.6 (Mistral 7B and Vicuna 13B) [26], LLaVA-Med [27], and CXR-LLaVA [28]. To ensure a fair comparison, these models, designed for 2D imaging tasks, are tested using Digitally Reconstructed Radiographs (DRRs) of 3D chest CT volumes—a method that ensures accurate representation in a 2D context, as suggested by [47], rather than using random or central 2D slices of a 3D volume (see benchmarking the CT-CHAT model section of Methods). In contrast, CT-CHAT directly processes 3D chest CT, giving it a significant advantage in understanding and interpreting the complex anatomical and pathological information within these scans.

### v4

@@ Abstract
Advancements in medical imaging AI, particularly in 3D imaging, have been limited due to the scarcity of comprehensive datasets. We introduce CT-RATE, a public dataset that pairs 3D medical images with corresponding textual reports. CT-RATE comprises 25,692 non-contrast 3D chest CT scans from 21,304 unique patients. Each scan is accompanied by its corresponding radiology report. Leveraging CT-RATE, we develop CT-CLIP, a CT-focused contrastive language-image pretraining framework designed for broad applications without the need for task-specific training. We demonstrate how CT-CLIP can be used in multi-abnormality detection and case retrieval, and outperforms state-of-the-art fully supervised models across all key metrics. By combining CT-CLIP’s vision encoder with a pretrained large language model, we create CT-CHAT, a vision-language foundational chat model for 3D chest CT volumes. Finetuned on over 2.7 million question-answer pairs derived from the CT-RATE dataset, CT-CHAT underscores the necessity for specialized methods in 3D medical imaging. Collectively, the open-source release of CT-RATE, CT-CLIP, and CT-CHAT not only addresses critical challenges in 3D medical imaging but also lays the groundwork for future innovations in medical AI and improved patient care.
Main text
@@
Three-dimensional (3D) medical imaging, such as computed tomography (CT) and magnetic resonance imaging (MRI), provides a more detailed and comprehensive view of a patient’s health status compared to two-dimensional (2D) imaging [1]. Among these, CT stands out due to its speed and widespread availability, making it the preferred modality for diagnosing various abnormalities [2]. As global demand for CT scans continues to grow, the need for rapid and accurate diagnostic results becomes more pressing. However, the manual generation of reports from these complex images is time-consuming and prone to errors, underscoring the urgent need for automation [3]. Artificial intelligence (AI) has the potential to address this challenge by efficiently identifying pathologies, reducing the workload on radiologists, and potentially saving lives in situations where immediate radiological expertise is not available.
@@
AI has already begun to revolutionize the interpretation of medical images [4, 5, 6, 7, 8]. Recent advances in automated diagnosis, classification, and segmentation have been driven by numerous publicly available datasets [9, 10, 11, 12, 13]. Beyond supervised methods, general AI systems have been developed using open-source image-text pairs in self-supervised learning, particularly in 2D imaging [14]. These versatile AI foundation models integrate comprehensive natural language annotations into the learning process without requiring task-specific training, thereby eliminating the need for explicit manual image annotations. Unconstrained by predefined classes or tasks during training, they enhance adaptability and have been successfully applied to tasks such as disease classification [15], case retrieval [16], and interactive diagnostics via chat interfaces [14]. Despite these advancements in 2D imaging, the application of advanced AI techniques to 3D imaging remains underexplored. A major challenge is the lack of open-source 3D medical imaging datasets paired with textual reports, which require extensive computational resources, specialized knowledge, and technical expertise to create [17, 18, 19]. Currently, the largest 3D medical imaging datasets available contain only a few thousand cases and do not include paired textual data [20, 21, 22], limiting the development of advanced AI models. Moreover, 3D models face higher computational demands and have fewer available pretrained open-source models compared to their 2D counterparts [23, 24]. Previous studies have only begun to address these challenges; for example, one involving chest CT is limited by its supervised approach, restricting detection to predefined abnormalities and offering only a small portion of its dataset publicly without paired textual reports [20].
@@
To address these challenges, we introduce CT-RATE, a dataset that pairs 50,188 reconstructed 3D chest CT volumes (over 14.3 million 2D slices) from 25,692 scans of 21,304 patients with corresponding radiology reports. Each volume is annotated with 18 abnormalities, supporting extensive validation and fine-tuning. CT-RATE is fully open-source. Leveraging CT-RATE, we develop CT-CLIP, a chest CT-focused contrastive language-image pretraining framework. CT-CLIP integrates extensive natural language annotations into the learning process, enabling the model to comprehend image-based semantic knowledge and perform various tasks without specific training. Our evaluation shows CT-CLIP’s ability to detect multiple abnormalities in a zero-shot setting [25], outperforming state-of-the-art fully supervised methods on three test sets from two different countries. We extend CT-CLIP’s capabilities by combining its vision encoder with a pretrained large language model to create CT-CHAT, a vision-language chat model for 3D chest CT volumes. CT-CHAT is fine-tuned with over 2.7 million question-answer pairs derived from the CT-RATE dataset. CT-CHAT outperforms open-source conversational vision-language AI assistants, such as LLaVA 1.6 [26], LLaVA-Med [27], and CXR-LLaVA [28], underscoring the need for specialized tools in 3D medical imaging. We also benchmark CT-CHAT against state-of-the-art 3D medical report generation methods, including RadFM [29] and CT2Rep [30], using clinically meaningful accuracy metrics. On two independent test sets, CT-CHAT outperforms these baselines by a significant margin, demonstrating both its clinical relevance and technical efficacy in report generation. In addition to quantitative benchmarks, we conduct expert evaluations with three board-certified radiologists, who consistently rated CT-CHAT’s reports as superior to those from RadFM and CT2Rep. While these results are encouraging, the experts emphasized that substantial progress is still needed for clinical deployment in 3D medical imaging. Our work marks a first step toward that goal, but important challenges remain.
@@ Results
We curate and open-source CT-RATE, a dataset comprising 25,692 non-contrast 3D chest CT volumes and corresponding radiology text reports from 21,304 unique patients (Figure 1a). These 3D volumes, reconstructed using various methods to suit different window settings [31], total over 14.3 million 2D slices. Each CT exam is accompanied by a report that includes an impression and a findings section, authored by a radiologist, along with additional sections detailing patient information and scan technique. CT-RATE is divided into training and internal validation sets for experiments (see creating the CT-RATE dataset section of Methods). While CT-RATE is a single-center dataset, it captures substantial heterogeneity in CT scanners, acquisition protocols, and reconstruction techniques.
To evaluate the generalization of our models (developed on CT-RATE) under distribution shifts, we incorporate two external validation sets from independent studies conducted in different countries [20, 32] (see adding external validation datasets to CT-RATE section of Methods). First, we use the public RAD-ChestCT dataset, which includes multi-abnormality labels, to assess model generalizability. Second, we perform a real-world validation by inviting an independent clinical research team (uninvolved in model development) to download our open-source codebase, pretrained models, and user interface from our public repository and apply them to their own local, private dataset without any adaptation. The performance scores they obtain are presented in the following sections. In particular, none of the external validation data is used during development; all model development is conducted solely on CT-RATE.
@@
Figure 2 summarizes CT-RATE alongside the open-source external validation set, RAD-ChestCT. CT-RATE spans a wide age range (18-102 years) (Figure 2a) and includes diverse pathologies primarily related to the heart and lung (Figure 2b). The reports vary in length and are authored by numerous radiologists, reflecting a variety of writing styles (Figure 2c). CT studies in CT-RATE also vary in slice numbers (Figure 2d), slice thickness, axial pixel (XY) spacing (Figure 2e), and include different manufacturers and a balanced sex distribution (Figure 2f). This heterogeneity is crucial for developing robust models capable of handling out-of-distribution cases. Supplementary Table 1 provides detailed statistics on abnormalities, manufacturers, and sex distributions in the training and internal validation sets. In conclusion, CT-RATE does not represent a curated study dataset but rather reflects real-world scenarios characterized by heterogeneity. This diversity ensures the models trained on CT-RATE are well-suited for deployment in clinical settings, offering reliable performance across a broad range of conditions.
@@
Unlike other methods using supervised machine learning for multi-abnormality classification of CT volumes, our study employs radiology reports, which provide a rich source of semantic knowledge. This significantly enhances the understanding of CT volumes and supports a wide range of potential zero-shot applications. Accordingly, we develop CT-CLIP, the first foundational model for chest CT-focused contrastive language-image pretraining, adapting a technique from a previous study [36]. We use both the impression and findings sections of radiology reports combined for CT-CLIP due to their superior performance compared to using either one alone, as demonstrated by the abnormality classification scores in an ablation study (Supplementary Figure 1a). While the best performance is achieved training on both impression and findings together, using only the findings section performs better than using only the impression section, indicating that the findings section contains more clinically informative content.
@@
We compare our zero-shot approach to this supervised baseline and observe significantly higher performance metrics for our approach. On the internal validation set from CT-RATE, the zero-shot CT-CLIP model achieves a mean AUROC that is 0.102 higher and a mean F1 score that is 0.050 higher than the supervised baseline. On the first external validation set, RAD-ChestCT, CT-CLIP outperforms the baseline with a mean AUROC gain of 0.085 and an F1 score improvement of 0.073. Similarly, on the second external validation set from UPMC, CT-CLIP maintains strong performance with a 0.087 increase in mean AUROC and a 0.037 boost in F1 score. Results on the external validation sets demonstrate that the zero-shot CT-CLIP model performs well not only on similar data distributions but also under distribution shifts, indicating its strong generalization to external datasets. A two-sided unpaired permutation test, using 10,000 permutations, highlights a statistically significant improvement $(p<0.05)$ in AUROC as a zero-shot classifier over the supervised (baseline) model, detailed in Supplementary Table 5.
Extended Data Figures 3, 4, 5 provide scores for all abnormalities on the CT-RATE internal, RAD-ChestCT external and UPMC external validation sets, respectively. Supplementary Figures 3, 7 show confusion matrices on the CT-RATE internal and RAD-ChestCT external validation sets, respectively. Notably, on the internal validation set, all abnormalities (18 out of 18) exhibit higher scores in our zero-shot approach compared to the supervised baseline in mean AUROC. Surprisingly, CT-CLIP excels by just leveraging existing radiology text reports in electronic health records, thus avoiding the need for expensive manual annotations. Moreover, CT-CLIP’s zero-shot capability offers flexible predictions beyond predefined classes and adapts efficiently to changing needs post-training.
Since the availability of sufficiently large datasets is essential for machine learning, especially for transformers, which require more data than convolutional neural networks [39, 40], we explore the impact of dataset size on classification performance. We train CT-CLIP on different portions of the CT-RATE dataset, specifically 9.8% (equivalent to the open-source portion of the RAD-ChestCT dataset [20], currently the largest publicly available chest CT dataset), 20%, 40%, 60%, 80%, and 100%. Figure 4c reports the classification performance across different dataset sizes. Notably, a consistent trend emerges: larger dataset sizes correlate with improved scores. This finding highlights the role of dataset size in enhancing CT-CLIP’s zero-shot classification performance and underscores the value of our extensive CT-RATE dataset for computational research. It also offers a clear avenue for future improvements, suggesting that further expanding the dataset or incorporating more diverse data could lead to even better performance.
@@
To enhance CT-CLIP’s detection performance, we explore a task-specific finetuning method, ClassFine, inspired by conventional linear probing, by adding a new classification layer to the pretrained vision transformer (Figure 3a). Supplementary Figure 1b indicates that training only the final layer may be sufficient for finetuning (see finetuning CT-CLIP for multi-abnormality detection section of Methods). While the ClassFine approach improves detection performance (Figure 4a), it restricts the model to predefined labels as any supervised classification method. Extended Data Figures 3, 4, 5 present abnormality scores, with confusion matrices in Supplementary Figures 5, 9 on the internal and external validation sets, respectively. To address the limitations of ClassFine, we develop the VocabFine approach, a novel open-vocabulary finetuning method inspired by WISE-FT [41] (see finetuning CT-CLIP for multi-abnormality detection section of Methods). VocabFine, like CT-CLIP’s zero-shot inference, is finetuned to improve label prediction accuracy (Figure 3c). Specifically, we use positive and negative prompts for abnormalities, calculate the logits with pretrained CT-CLIP, and align these logits with the ground truth labels. While the finetuned model with VocabFine retains the open-vocabulary inference capability of CT-CLIP (Figure 3d), it also consistently outperforms ClassFine (Figure 4a), further emphasizing the value of our open vocabulary finetuning. Abnormality-specific improvements are detailed in Extended Data Figures 3, 4, 5 while Supplementary Figure 1b underscores the importance of training all weights in finetuning. Supplementary Figures 4, 8 provide confusion matrices on the internal and external validation sets, respectively. Notably, both ClassFine and VocabFine are trained solely on CT-RATE; thus, their generalization to external sets (RAD-ChestCT and UPMC) demonstrates the robustness of CT-CLIP’s pretrained representations under distribution shifts. This performance advantage across multiple test sets supports the effectiveness of our finetuning strategies. However, despite these promising results, the abnormality detection performance still falls short of clinical deployment thresholds.
@@
Beyond its superior performance, VocabFine retains the open-vocabulary capabilities of the original CT-CLIP model, enabling inference on abnormalities that were not explicitly included during fine-tuning. This makes the model inherently more flexible and clinically useful in real-world scenarios where emerging or rare findings may need to be recognized without retraining. Additionally, unlike ClassFine, which only adds a classification head on top of the vision encoder and is therefore restricted to vision-only downstream tasks, VocabFine preserves and fine-tunes both the vision and text encoders. This dual-modality design allows it to generalize to a wider range of tasks, such as text-to-image retrieval and multimodal question answering, which rely on aligned visual and linguistic representations. In summary, our VocabFine approach improves multi-abnormality detection accuracy and also maintains the foundational model’s versatility with adaptability across diverse downstream applications.
@@
Beyond zero-shot detection, CT-CLIP serves as a powerful tool for image-to-image retrieval [42], identifying the relevant volumes from a query volume by comparing cosine similarities between the embeddings (Figure 5a). We test on both CT-RATE internal and RAD-ChestCT external validation sets using the vision transformer of CT-CLIP, finetuned models, and the supervised baseline. We use the Mean Average Precision at K (MAP@K) metric, where K represents the number of returned 3D chest CT images (see benchmarking the CT-CLIP models for CT volume retrieval section of Methods). As K increases, MAP@K scores typically decrease due to less accurate matches. Results, as in Figure 5c, show comparable performance among methods, with all significantly outperforming random retrieval. Although the external validation set from RAD-ChestCT has higher MAP@K scores, the fold increase over random retrieval is lower (three times on the external vs. six times on the internal set), due to differing abnormality distributions. This suggests that examining fold change offers a more accurate comparison. Beside, finetuning models does not significantly alter the vision transformer’s weights, as CT-CLIP and the finetuned models perform similarly across MAP@K metrics.
@@
CT-CHAT’s architecture leverages the strengths of CT-CLIP by using its pretrained vision transformer to encode 3D chest CT volumes. These encodings are then passed to a large language model (LLM) through a multimodal projector, as shown in Figure 6a. We perform an ablation study evaluating CT-CHAT’s performance with 4 different state-of-the-art LLMs, as detailed in Extended Data Figure 6. We select 70B Llama 3.1 [44], outperforming others in key metrics, for the optimal configuration of CT-CHAT.
@@
Given the absence of generalist models for 3D imaging, we evaluate CT-CHAT against several state-of-the-art open-source vision-language AI models, including LLaVA 1.6 (Mistral 7B and Vicuna 13B) [26], LLaVA-Med [27], and CXR-LLaVA [28]. To ensure a fair comparison, these models, designed for 2D imaging tasks, are tested using Digitally Reconstructed Radiographs (DRRs) of 3D chest CT volumes, a method that ensures accurate representation in a 2D context, as suggested by [46], rather than using random or central 2D slices of a 3D volume (see benchmarking the CT-CHAT model section of Methods). In contrast, CT-CHAT directly processes 3D chest CT, giving it a significant advantage in understanding and interpreting the complex anatomical and pathological information within these scans.